Pet Lodge Inc., which offers day and overnight kennel services for pets, is offering a promotion for new customers signing a contract: purchase 3 overnight stays, a 10-visit day pass punch card, and a new leash for $182. The total transaction price of $182 is stated to the customer as $63 for the overnight stay, $105 for the punch card, and $14 for the leash. Allocate the $182 transaction price under the following separate scenarios.

Note: When answering the following questions, round the allocated transaction price to the nearest dollar.

a. The three items (overnight stays, day pass punch card, and leash) are sold separately by Pet Lodge Inc. The standalone selling prices of the overnight stays, day pass punch card, and a new leash are $63, $105, and $14, respectively.

b. The three items (overnight stays, day pass punch card, and leash) are sold separately by Pet Lodge Inc. The standalone selling prices of the overnight stays, day pass punch card, and leash are $75, $126, and $18, respectively

c. Two of the three items (overnight stay and day pass punch card) are sold separately by Pet Lodge Inc. The standalone selling prices of the overnight stays and day pass punch card are $63 and $105, respectively. Because Pet Lodge Inc. has never sold leashes prior to this promotion, Pet Lodge Inc. determined the average selling price of similar leashes in the market place to be $18.

d. Two of the three items (overnight stays and day pass punch card) are sold separately by Pet Lodge Inc. The standalone selling prices of the overnight stays and day pass punch card are $63 and $105, respectively. Pet Lodge Inc. purchased the leashes for $11 each and expects a gross profit margin of 50%.

e. Two of the three items (overnight stays and day pass punch card) are sold separately by Pet Lodge Inc. The standalone selling prices of the overnight stays and day pass punch card are $63 and $105, respectively. Pet Lodge Inc. has never sold leashes prior to this promotion, and is not able to obtain a reliable comparable market price because the leashes are unique and there is nothing comparable on the market.
Answer #1

Given Transaction price is $182.

Performance obligations are separately identifiable and they are

  • Overnight Stays
  • Day Pass Puncha Card
  • New Leash

Allocation of Transaction price to separate scenerios:

a)

The Combined price of standalone selling prices = 63+105+14

= 182

As combined price is equal to transaction price, same prices are the standalone selling prices from the transaction.

The Combined price of standalone selling prices = 63 +105+14

= 182

Standalone Selling Price from Transaction = (SSP/Combined price)*Transaction price

Overnight Stays

= (63/182)*182

= $63

Day Pass Punch Card

= (105/182)*182

= $105

New Leash

= (14/182)*182

= $14

b)

The Combined price of standalone selling prices = 75 +126+18

= 219

Standalone Selling Price from Transaction = (SSP/Combined price)*Transaction price

Overnight Stays

= (75/219)*182

= $62

Day Pass Punch Card

= (126/219)*182

= $105

New Leash

= (18/219)*182

= $15

c)

As SSP for new leash is not available and Pet Lodge Inc never sold it before but market price is available, we can assume $18(Given) as market assessment price for SSP of new leash.

The Combined price of standalone selling prices = 63 +105+18

= 186

Standalone Selling Price from Transaction = (SSP/Combined price)*Transaction price

Overnight Stays

= (63/186)*182

= $62

Day Pass Punch Card

= (105/186)*182

= $103

New Leash

= (18/186)*182

= $17 (Can be 18 also but rounded off to 17 to match the total transaction price)

d)

As SSP for new leash is based on Cost plus Margin approach, we need to calculate SSP for New Leash.

= $11 + 50% of gross profits

(Sales-COGS)/Sales = 50%

Let we assume sales be x

(x-11)/x = 50%

x-11 = 0.5x

0.5x = 11

x = 11/0.5

= 22

SSP of New Leash = $22

The Combined price of standalone selling prices = 63 +105+22

= 190

Standalone Selling Price from Transaction = (SSP/Combined price)*Transaction price

Overnight Stays

= (63/190)*182

= $60

Day Pass Punch Card

= (105/190)*182

= $101

New Leash

= (22/190)*182

= $21

e)

As SSP for new leash is not available.

  • It is not sold before
  • No reliable market price available

In this case, we need to use residual approach where SSP of unknown would be the residual value after deducting known SSPs from transaction price.

Transaction Price $182

Less: SSP of Overnight Stay ($63)

Less: SSP of Punch Card    ($105)

SSP of new leash = 182 - 168 = $14

As combined price is equal to transaction price, same prices are the standalone selling prices from the transaction.

The Combined price of standalone selling prices = 63 +105+14

= 182

Standalone Selling Price from Transaction = (SSP/Combined price)*Transaction price

Overnight Stays

= (63/182)*182

= $63

Day Pass Punch Card

= (105/182)*182

= $105

New Leash

= (14/182)*182

= $14